### v1

# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

Abstract
Image–Language Foundation Models (ILFM) have demonstrated remarkable success in image–text understanding/generation tasks, providing transferable multimodal representations that generalize across diverse downstream image-based tasks. The advancement of video-text research has spurred growing interest in extending image-based models to the video domain. This paradigm, known as image-to-video transfer learning, succeeds in alleviating the substantial data and computational requirements associated with training video-language foundation models from scratch for video-text learning. This survey provides the first comprehensive review of this emerging field, which begins by summarizing the widely used ILFM and their capabilities. We then systematically classify existing image-to-video transfer learning strategies into two categories: frozen features and modified features, depending on whether the original representations from ILFM are preserved or undergo modifications. Building upon the task-specific nature of image-to-video transfer, this survey methodically elaborates these strategies and details their applications across a spectrum of video-text learning tasks, ranging from fine-grained (e.g., spatio-temporal video grounding) to coarse-grained (e.g., video question answering). We further present a detailed experimental analysis to investigate the efficacy of different image-to-video transfer learning paradigms on a range of downstream video understanding tasks. Finally, we identify prevailing challenges and highlight promising directions for future research. By offering a comprehensive and structured overview, this survey aims to establish a structured roadmap for advancing video-text learning based on existing ILFM, and to inspire future research directions in this rapidly evolving domain.
Index Terms:
Foundation Models, Video Understanding, Transfer Learning, Vision-Language, Image-to-Video Adaption.
I Introduction
Multimodal learning, particularly vision-language (or vision-text) learning, has attracted significant attention in the community in recent years. With larger datasets and more powerful computing resources, researchers can train image-language foundation models that can adapt to many downstream tasks and modalities. Image-based multi-modality systems such as CLIP[1] and BLIP[2] showed that training models with pairs of images and texts at scale can yield widely transferable features. Newer models like MDETR[3] and GroundingDINO[4] go further by aligning free-form language with locally aligned image regions, closing the gap between textual description and image location. Building on these successes, researchers are actively exploring how to migrate these models, proven successful on static images, to the more complex video domain. The goal is to address the unique challenges posed by the additional temporal dimension in videos and to leverage the rich dynamic information.
Figure 1: Comparison of two representative strategies for addressing video-text understanding problem with image-text and video-text foundation models. Pretraining video-based foundation models is much more challenging than pretraining image-based foundation models, as they often contain more parameters and require larger amounts of training data and computational resources.
Over recent years, image-to-video transfer learning, which adapts knowledge from vision-language foundation models to the video domain, has emerged as a prominent paradigm for video-language understanding. Alternatively, another approach involves pre-training video-language foundation models directly from massive video-text data before transferring them to downstream tasks. As shown in Fig 1, compared to direct pre-training—which faces challenges such as high data collection costs and low computational efficiency—we find that initializing from large-scale image-text pre-trained models offers greater practicality and scalability. This is attributed to the greater abundance of image-text data and lower associated costs. For instance, UniFormerV1 [5] was pre-trained directly on the Kinetics-400 [6] dataset, whereas UniFormerV2 [7] underwent post-pre-training starting from an image-text model. Compared to V1, UniFormerV2 achieves superior performance with reduced training overhead, demonstrating that adapting image representations aligned with text to various video scenarios via task-specific transfer techniques is both more flexible and cost-effective.
However, extending the notable success of image-language foundation models to the video domain presents significant challenges, primarily arising from the temporal dimension. In contrast to static images, videos comprise sequences of frames that encode complex spatiotemporal dynamics, including motion, event progression, and causal interactions. Consequently, a fundamental architectural paradigm shift is necessary, moving from processing static spatial information to effectively modeling these intricate spatial-temporal relationships. Early methods processed each frame with 2D image networks and then applied a simple temporal fusion layer, but this setup cannot capture fine motion cues. Recently, researchers are adapting these models to video domain by inserting lightweight temporal modules into a frozen backbone and applying parameter-efficient fine-tuning techniques like adapters, often employing these strategies in hybrid configurations. Furthermore, a growing number of diverse approaches are continuously being introduced and explored in this domain. Side-tuning approaches [8, 9], for example, is becoming increasingly popular as one of the most promising ways to accommodate the critical challenge of image-to-video transfer learning brought by the huge parameter and computation overheads. Other methods like distillation, also shows further potential to adapt the image-text aligned knowledge into the video domain especially under the era of Multimodal Large Language Models (MLLMs) in which we can capitalize on the exceptionally strong multimodal knowledge contained by these excessively large models.
Despite the existence of numerous paradigms for image-to-video transfer learning, it remains underexplored which paradigm or combination thereof is most suitable for specific downstream tasks, as different tasks require models to possess distinct comprehension capabilities. For example, spatio-temporal video grounding (STVG) requires a model to enhance its capability for fine-grained temporal event localization while preserving its existing static spatial grounding performance. This objective is well-suited for the integration of supplementary video-level features. In contrast, Video Question Answering (Video QA) necessitates only a coarse-grained comprehension of video events but demands a more robust capacity for multi-modal knowledge understanding. The optimal transfer learning paradigm for this latter task remains an open question.
In this survey, we provide a comprehensive overview of established strategies for migrating image foundation models to various video-centric tasks. Beyond summarizing these techniques, we offer a critical analysis of their suitability for different application scenarios, concluding with a discussion on the prevailing challenges and promising directions for future research. The rest of this survey is organized as follows. Section II introduce the commonly used vision-language foundation models developed in the community. Section III reviews and categorizes the techniques of leveraging image-level knowledge from Image-Language Foundation Models (ILFM) for video-language understanding. Section IV and V systemically review the image-to-video transfer learning approaches for addressing varied downstream video-text understanding tasks including fine-grained to coarse-grained problems, respectively. In Section VI, we provide the detailed experimental analysis. Finally, we provide several promising image-to-video transfer learning research directions in Section VII and conclude in Section VIII.
II Overview of Vision-Language Foundation Model
Pretrained ILFM that leverage both visual and textual information have revolutionized many visual perception tasks including object detection [10, 11, 12], image captioning [13, 14, 15], and visual question answering (VQA)[16, 17, 18], etc. Among the most influential models in this domain, several have introduced foundational innovations: CLIP[1] establishes image-text alignment through contrastive pretraining; BLIP[2] and BLIP-2[19] improve vision-language representation learning with bootstrapped captioning and parameter-efficient architecture; MDETR[3] and GroundingDINO[4] advance open-vocabulary detection and referring comprehension using language-aware visual encoding; and LLaVA[20] integrates visual encoders with large language models to support instruction-based multimodal dialogue. Collectively, these models bridge the gap between vision and language understanding and provide a diverse set of powerful and adaptable foundations for a wide spectrum of downstream vision-language applications.
TABLE I: Classification and Analysis of Image-Language Foundational Models
Foundation Model Core Architecture Pre-training Objective(s) Pre-training Data Primary Output Key Capability
CLIP[1] Dual-Encoder (Image Encoder + Text Encoder) Image-Text Contrastive Learning 400M image-text pairs (Web-crawled) Aligned Image & Text Embeddings Zero-Shot Image Classification, Robust Feature Representation
BLIP & BLIP2[2, 19] Multimodal Mixture of Encoder-Decoder (ViT, Q-Former, Text Encoder/Decoder) Contrastive, Matching, Captioning Approx. 14M image-text pairs (Academic datasets + Web data) Embeddings, Matching Scores, Text Visual Question Answering (VQA), Image Captioning
MDETR[3] Transformer Encoder-Decoder Phrase-to-Box Grounding 1.3M images with 11.5M phrase-box annotations (Flickr30k, GQA, etc.) Bounding Boxes, Class Labels Open-Vocabulary Object Detection & Grounding
GroundingDINO[4] Transformer-based Detector with Image & Text Backbones Open-Set Object Detection Combination of COCO, Objects365, GoldG, etc. (Detection & Grounding datasets) Bounding Boxes, Class Labels Open-Set Detection
LLaVA[20] Vision-Language Model(Image Encoder + Projection + LLM) Instruction Tuning on Image-Text Data Approx. 158K language-image instruction-following samples (GPT-4 generated) Textual Response Conversational VQA, Visual Reasoning
MDETR111ghttps://github.com/ashkamath/mdetr [3] is a foundational vision-language model designed for grounded visual-language understanding, where it predicts object locations in conjunction with their corresponding textual descriptions. As illustrated in Figure 2(a), the model employs dual encoders: a convolutional visual encoder (e.g., ResNet-101[21] or EfficientNet[22]) and a transformer-based text encoder (e.g., RoBERTa[23]). These encoders project both images and text into a shared latent space, where semantic alignment is enforced by maximizing similarity for matched pairs and minimizing it for mismatched ones. This joint representation allows MDETR to reason over spatial and semantic relationships across modalities, making it particularly effective for tasks such as phrase grounding, referring expression segmentation, and visual question answering (VQA).
MDETR is pre-trained on a curated corpus of approximately 200k images encompassing around 1.3 million region–phrase pairs.These aligned pairs are sourced from diverse datasets including Flickr30k[24], MS-COCO[25], and Visual Genome[26], ensuring dense grounding of objects to textual descriptions. The model employs a graph-coloring technique to consolidate multiple phrases per image, ensuring non-overlapping and semantically coherent region assignments that facilitate more effective representation learning. Trained at this scale, MDETR demonstrates strong zero-shot generalization capabilities across a variety of downstream tasks, including open-domain visual retrieval and fine-grained visual recognition.
GroundingDINO [^2] https://github.com/IDEA-Research/GroundingDINO[4] introduces a novel approach to open-set visual grounding, designed to localize and identify objects based on free-form textual descriptions. As illustrated in Figure 2(b), the model builds upon a transformer-based architecture that integrates visual and linguistic modalities through a dedicated fusion module, enabling fine-grained cross-modal alignment between image regions and text queries. This design allows the model to effectively ground natural language phrases in images, making it highly suitable for tasks such as referring expression comprehension and open-vocabulary detection.The model is trained on a large-scale multi-source dataset comprising image-text pairs with bounding box annotations. Key data sources include Flickr30K Entities [24], MS-COCO [25], Visual Genome [26], and Objects365 [27], which provide diverse and fine-grained grounding supervision across common and long-tail categories. This extensive pretraining enables robust representation learning for both visual and textual modalities.
A major innovation of GroundingDINO lies in its integration of the DINO (Distilled Knowledge) framework into visual grounding. Through cross-modal knowledge distillation, the model captures fine-grained semantic information from textual descriptions and transfers this understanding to visual contexts with high precision. This mechanism enables GroundingDINO to localize objects in unseen images using only text prompts—without any task-specific retraining—showcasing remarkable zero-shot capability. This strength makes it particularly suitable for applications in dynamic visual environments and open-ended image retrieval tasks.
Figure 2: Popular image-textual foundation models developed in the community, which are widely used in the image-to-video transfer learning for addressing downstream video-language understanding problems.
CLIP (Contrastive Language-Image Pre-training) [^3] https://github.com/openai/CLIP [1] is an image-language foundational model that learns joint representations of images and text through contrastive pre-training. As Figure 2(c) illustrate, its architecture consists of dual encoders: a vision encoder built on scalable neural networks such as vision transformers or convolutional architectures, and a text encoder based on transformer networks. These encoders are trained to map paired images and texts into a shared latent space, where semantic alignment is achieved by maximizing similarity scores for correct image-text pairs and minimizing them for incorrect ones. The model is pre-trained on a dataset of 400 million publicly available image-text pairs collected from the web, enabling broad semantic understanding. A key innovation of CLIP is its zero-shot generalization capability, enabling it to be transferrable to other perception tasks, such as image classification and cross-modal retrieval, without requiring task-specific fine-tuning. Empirical studies show that CLIP performs competitively with fully supervised models across 30 diverse vision benchmarks, particularly excelling in fine-grained recognition and open-domain retrieval.
Image-to-Video Transfer Learning
   Frozen Features (§III-A)    Knowledge Distillation       (§III-A1) Fine-grained: OVTrack [28],OVTrack+ [29],VOVTrack [30],SLAck [31],GLATrack [32],UniVTG [33]. Coarse-grained: VideoDistill [34], Clip2tv [35], Text4Vis [36], FitCLIP [37], CLIPPING [38]. TeachCLIP [39],    Post-Network Tuning       (§III-A2) Fine-grained: NetTrack [40], OVTR [41], LaMOTer [42], OV2Seg [43], UnLoc [33], CLIP-VIS [44], MomentDETR [45], BAM-DETR [46], UniVTG [33], UnLoc [47], TFVTG [48], CoSPaL [49], STPro [50], BriVIS [51], OVFormer [52], DVIS++ [53]. Coarse-grained: BIMBA [54], LinVT [55], CLIP4Caption [56], E2DVC [57], HierarQ [58], UniFormerV2 [7].       Side-Tuning       (§III-A3) Fine-grained: R2-tuning [59], VDI [60], PATF [61], Tune-A-Video [62]. Coarse-grained: CLIP2Video [63], STAN [64], EVL [65], DiST [66], MAMS [67]. Modified Features (§III-B)       Full Fine-Tuning       (§III-B1) Fine-grained: STVGBERT [68], SWINBERT [69], TubeDETR [70], STCAT [71], ReferDINO [72], GEN1 [73]. Coarse-grained: Stable Video Diffusion [74], Stable Diffusion [75], I2VGen-XL [76], CLIP4Clip [77], CLIP-Hitchhiker [78], ViFi-CLIP [79], TVR [80], CenterCLIP [81].       Partial Tuning       (§III-B2) Fine-grained: Tune-A-Video [62], FRESCO [82]. Coarse-grained: VideoCrafter1 [83], TC-LLaVa [84], ModelScope Text-to-Video [85], TEA [86], mPLUG-2 [87]. Fine-Tuning with Extra Models       (§III-B3) Fine-grained: iKUN [88], CoSD [89], CoSTA [90], CGSTVG [91], TASTVG [92], Vtimellm [93], Seq2Time [94], NumPro [95], Chrono [96], NetTrack [40], InstFormer [97], Grounded-SAM-2 [98], AL-Ref-SAM-2 [99]. Coarse-grained: VTG-LLM [100], X-pool [101], MoMa [102], Vid2Seq [103], CM2 [104], EvCap [105], BIKE [106].    Fine-Tuning with Adapter       (§III-B4) Fine-grained: MASA [107], LlaViLo [108], ReVisionLLM [109], iKUN [88]. Coarse-grained: Tem-Adapter [110], ST-Adapter [111], Adaptformer [112], AIM [113], DUALPATH [114], TS2-Net [115], X-CLIP [116], D2ST [117], TR-Adapter [118].     Fine-Tuning with LoRA       (§III-B5) Fine-grained: ReferDINO [72], VTG-LLM [100], VTimeLLM [93], ReVisionLLM [109], NumPro [95], InstFormer [97]. Coarse-grained: DiscoVLA [119], HierarQ [58].       Prompt Tuning       (§III-B6) Coarse-grained: ATP [120], ActionCLIP [121], X-CLIP [116], ViTiS [122], Q-ViD [123], X-CLIP [124]. VOP [125], Clip-vip [126].
Figure 3: Taxonomy of Image-to-video Transfer Learning Methods. Existing methods are categorized according to the employed transfer learning mechanisms, forming two major groups: frozen features and modified features. The frozen-feature group includes approaches such as knowledge distillation, post-network tuning, and side-tuning, while the modified-feature group encompasses various fine-tuning strategies, including full fine-tuning, adapter-based tuning, LoRA, and prompt tuning. Each category is further divided into fine-grained and coarse-grained paradigms, reflecting different levels of temporal and semantic alignment between visual and textual modalities.
BLIP [^4] https://github.com/salesforce/BLIP [2] (Bootstrapping Language-Image Pre-training) is a unified vision-language framework designed to jointly optimize understanding and generation tasks through multimodal co-training. As Figure 2(d) shows, its architecture integrates a vision encoder for image feature extraction, a text encoder-decoder for language processing, and a cross-modal transformer that dynamically fuses visual and textual representations. This hybrid design enables versatile applications, including visual question answering, image captioning, and text-guided image retrieval.
BLIP is effectively trained with a two-stage pre-training strategy. In the first stage, it employees a data bootstrapping mechanism to filter out noisy web-collected image-text pairs and generate high-quality synthetic captions via knowledge distillation. In the second stage, it optimizes the parameters by considering image-text contrastive learning to align modalities, image-text matching to verify cross-modal consistency, and generative tasks to produce contextually relevant textual outputs. Pre-trained on a combination of 129 million curated and 140 million web-sourced image-text pairs, BLIP achieves state-of-the-art results across 14 benchmarks, demonstrating strong robustness in low-data regimes.
BLIP-2555https://github.com/salesforce/LAVIS/tree/main/projects/blip2 [19] further improves BLIP by additionally introducing a lightweight querying module to bridge pre-trained vision encoders and large language models, in order to align visual features with textual embeddings. This delivers significant accuracy gains on some downstream visual-textual understanding tasks like visual question answering, with introducing considerable computational overhead. BLIP and its extensions BLIP-2 have been widely used in the community for addressing multimodal reasoning tasks including healthcare diagnostics and interactive recommendation systems.
LLaVA [^6] https://github.com/haotian-liu/LLaVA (Large Language and Vision Assistant) [20] is a vision-language model that integrates visual understanding with natural language processing through an end-to-end trainable architecture. As Figure 2(e) shows, The model consists of a pre-trained vision encoder CLIP-ViT-L/14 and a large language model (LLM) to process both textual and visual features. LLaVA employs a simple projection layer to align visual representations with the language model’s embedding space, enabling cross-modality fusion.
The training of LLaVA is split into two stages to ensure better cross-modality alignment and task-specific adaptability. In the first stage (i.e., feature alignment pretraining), the projection layers in LLaVA are learned with freezing vision encoder, which aims to learn a good semantic alignment between visual and textual representations. The pretraining is conducted on a subset of CC3M (Conceptual Captions 3M)[127], a dataset comprising 3 million image-text pairs. In the second stage (instruction fine-tuning), a large scale of instruction-following data generated by GPT-4, which includes approximately 158K multi-modal instruction-response pairs covering diverse tasks such as visual dialogue, detailed descriptions, and complex question answering are used to enhance the model’s reasoning and conversational capabilities. In this stage, both the projection layer and the LLM are updated and the vision encoder is fixed.
LLaVA-1.5 [128] extends LLaVA by replacing its image encoder with CLIP-ViT-L-336px and vision-language projection layer with MLPs, which achieve better multi-modality understanding. Moreover, the model also introduces a multi-patch partitioning strategy, which improves the efficiency of processing high-resolution images and facilitates dynamic-resolution inputs. It also extends the training data with academic benchmarks of varied vision tasks including visual question answer (e.g., OKVQA [129] and A-OKVQA [130]) and OCR-intensive tasks (e.g., OCRVQA[131] and TextCaps [132]). LLaVA-NeXT [133] improves LLaVA-1.5 by quadrupling input image resolution (up to 672×672) and supporting flexible aspect ratios inputs, enabling capturing finer visual detail. It enhances reasoning, OCR, and world knowledge through optimized visual instruction tuning while maintaining data efficiency with under 1M training samples. The model achieves superior performance across benchmarks, even surpassing Gemini Pro [^7] https://deepmind.google/models/gemini/pro/ in some tasks, while retaining LLaVA’s design and fast training scalability (e.g., 34B parameters trainable in nearly 1 day on 32 A100 GPUs).
Discussion: All these foundation models not only provide a solid foundation for improving image understanding in the presence of natural language but also set the stage for their application in more complex video analysis tasks, where the integration of dynamic temporal information becomes essential. Among these image-based foundation models, MDETR and GroundingDINO demonstrate stronger capabilities in capturing fine-grained visual–textual correspondences, as they are explicitly trained with grounding supervision to align object-level regions with textual phrases. By contrast, models such as CLIP and BLIP are more adept at handling abstract or high-level visual–textual correspondences, largely due to large-scale contrastive or generative pretraining. Instruction-tuned models like LLaVA bridge the two extremes by leveraging large language models to reason over both detailed and abstract alignments, though often at the expense of precision in object-level grounding.
III Overview of Image-to-Video Transfer Learning
In this section, we summarize common transfer paradigms for adapting image-level knowledge from Image-Language Foundation Models (ILFM) to address video-text learning tasks. Depending on whether the vision-language features extracted by ILFM remain unchanged or are adapted, we categorize existing transfer learning methods into frozen features and modified features, which will be elaborated in the following subsections.
III-A Frozen Features
Figure 4: Methods and architectures of transferring pre-trained image-text model to video domain via frozen features.
By treating video as sequence of image frames in temporal order, many existing methods directly leverage pre-trained ILFM to extract frame-wise cross-modal features and then use the frozen frame-level features to facilitate their image-to-video transfer learning. Representative lines of work include knowledge distillation, post-network tuning, and side-tuning.
III-A1 Knowledge distillation
As shown in Figure 4(A), within the knowledge distillation strategy, ILFM is primarily employed as a teacher model to enhance the cross-modal alignment capabilities of image/text encoders in existing video-text understanding models. Various distillation objectives can be defined to facilitate targeted knowledge transfer across encoders, working in conjunction with the primary task loss. Although this paradigm has been widely adopted in early Referring Video Object Segmentation (RVOS) and Open-Vocabulary Multi-Object Tracking (OV-MOT) methods to introduce open-vocabulary capabilities from models like CLIP, it often overlooks the discrepancy between image-level and video-level representations. This oversight may introduce erroneous knowledge from ILFMs, consequently limiting the effectiveness of cross-modal video-text understanding.
III-A2 Post Network Tuning
Many methods achieve temporal modeling by additionally introducing a post network to process the frozen features obtained for each selected frame (shown in Figure 4(B)). It is straightforward to construct the post network with temporal attention mechanism, enabling the information exchange among varied frames. However, this paradigm struggles to effectively capture dynamic temporal patterns from static features, resulting in limited performance. To more effectively harness the visual potential of frozen large-scale image models, researchers employ a Q-Former [19] architecture, utilizing a set of learnable queries to interact with features from all video frames. Through this approach, queries can dynamically capture temporal dependencies and motion patterns in videos, rather than augmenting image features with temporal relations. This not only reduces computational costs but also unlocks the potential of ILFM. Beyond general temporal modeling, many methods have incorporated more task-specific temporal modeling strategies, such as the Hungarian algorithm or Brownian bridge-based association. In summary, this paradigm relies on high-quality image-text features and carefully designed, task-specific temporal modeling architectures to achieve strong video understanding performance.
III-A3 Side-Tuning
as illustrated in Figure 4(C), Side-tuning [8] introduces a learnable side network alongside the frozen image and text encoders of ILFM. It receives the layer-wise latent features extracted by ILFM encoders as input, facilitating hierarchical spatial modeling and temporal refinement in a coarse-to-fine manner. In practice, the side network adopts a lightweight design, incorporating both text-modulated spatial pooling and temporal modeling operators [59]. During training, gradients are confined within the side-blocks, ensuring both parameter and memory efficiency. Recent study also show that the system can be further enhanced by connecting side-blocks with the outputs of CLIP, which achieves better results on several downstream video-text understanding tasks [59].
Discussion: Frozen-feature-based image-to-video transfer learning frameworks offer notable advantages in adapting image-based foundation models to video–language understanding tasks. They maintain high parameter efficiency by freezing pretrained architecture and weights, thereby mitigating catastrophic forgetting of previously acquired knowledge. However, their performance may be limited by the inherent representational capacity of static image-level features and their imperfect alignment with video-specific task requirements, particularly when generalizing to novel scenarios.
III-B Modified Features
Instead of directly using foundation models to extract frozen features, many approaches perform image-to-video transfer learning by explicitly adjusting the backbone model architecture and parameters, which we classify as modified features. In the following, we will elaborate on the techniques involved in modified features including fully-finetune, finetune with extra models, finetune with partial tuning, finetune with adapter, finetune with LoRA, and finetune with prompt tuning, whose architectures are presented in Figure 5, respectively.
Figure 5: Methods and architectures of transferring pre-trained image-text model to video domain via modified features.
III-B1 Full Fine-Tuning
It is straightforward to empower a pre-trained image-text foundation model with the ability of video-text understanding by building a temporal module on the top of each block of image-language foundation model and then simply finetuning all the parameters involved in the temporal modules together with the ILFM [68, 70, 71, 69]. Although simple and effective, these approaches indeed are less applicable in practice as they need a large-scale well-defined and annotated data and a large amount of computing power in the model training stage.
III-B2 Partial Tuning
Partial tuning strikes a balance between full fine-tuning and more parameter-efficient methods like adapters or prompt tuning. This strategy involves updating only a specific subset of the pre-trained ILFM’s parameters, while freezing the rest of the model parameters. Recently, several methods [134, 135, 136] have introduced temporal attention modules into the original spatial attention modules, enabling ILFMs to acquire temporal–spatial modeling capability. Such tuning strategies strike a favorable balance between performance and computational cost, while still relying heavily on the inherent spatial capacity of the ILFM, as they do not involve full model training.
III-B3 Fine-Tuning with Extra Models
Owing to the limited capacity of image-level features in capturing dynamic information, numerous image-to-video transfer learning methods incorporate auxiliary models to enhance temporal modeling. Currently, two dominant paradigms exist. The first paradigm augments image-based features with dynamic signals (e.g., optical flow, 3D spatio-temporal features) and introduces specialized modules to fuse or align them. For example, in STVG, recent methods employ both image and video backbones (e.g., SlowFast[137] or Video Swin Transformer[138]), coupled with attention-based fusion modules, for collaborative spatio-temporal modeling. The second paradigm leverages the structural outputs of external models—such as detection results, segmentation masks, and motion saliency maps—to guide image-to-video transfer with lower computational cost. In summary, fine-tuning with extra models offers an effective strategy for injecting video-specific structural information into ILFMs. It is particularly suitable when video annotations are limited or noisy, and has demonstrated strong empirical performance across tasks such as referring video object segmentation, temporal grounding, and video-language retrieval.
III-B4 Fine-Tuning with Adapter
Adapter-based methods offer a parameter-efficient solution to adapting ILFMs for addressing video-language understanding tasks, which intends to insert a set of lightweight, trainable module (referred as adapters) into some intermediate layers of pre-trained ILFMs  [139, 140, 60]. The adapter module are typically defined as a stack of temporal attention and cross-frame interaction operators in practice. By keeping the parameters corresponding to ILFM frozen, adapter-based approaches can retain the ability of visual-semantic understanding gained in pre-trained ILFM while efficiently learn temporal patterns specific to video data through the newly introduced adapter. Recently, the adapter-based finetuning techniques are particularly attractive in scenarios with limited video annotations or computational budgets, and they are emerging as a prominent direction for addressing downstream tasks like video question answering, temporal video grounding, and video captioning.
III-B5 Fine-Tuning with LoRA
As a representative parameter-efficient fine-tuning (PEFT) method, Low-Rank Adaptation (LoRA) [141] provides an alternative paradigm for ILFM adapting. Instead of directly learning the updating of model weights $\Delta W$, LoRA select to approximate the updated weight with two small, trainable low-rank matrices, i.e., $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During fine-tuning, the original parameters of the ILFM remain frozen and locked, preserving the extensive knowledge acquired during pre-training. Only the injected low-rank matrices $A$ and $B$ are updated, which dramatically reduces the number of trainable parameters and the accompanying GPU memory footprint. This approach implicitly constrains the model’s exploration within a low-dimensional subspace, effectively mitigating the risk of overfitting, which is particularly beneficial when labeled video data is scarce. Consequently, LoRA has gained significant traction for efficiently adapting large-scale foundation models to various downstream video understanding scenarios, establishing itself as a cornerstone technique in resource-constrained environments. However, as LoRA itself is incapable of modeling temporal information, this paradigm relies on external temporal modeling modules or networks, such as those used in fully finetuning or post-network tuning.
III-B6 Prompt Tuning
Prompt tuning has recently emerged as a highly parameter-efficient alternative for adapting ILFMs to video-related tasks [122, 123]. By freezing the backbone and introducing a small set of learnable prompts, the model is guided to capture temporal cues without modifying its core parameters. A common design is to share a single set of prompts across all frames to enforce temporal consistency, or to propagate prompts sequentially across frames so that information accumulates and evolves with motion. These temporal variants enable the frozen ILFM to move beyond static frame modeling and capture motion cues and cross-frame dependencies. This strategy achieves efficient spatio-temporal adaptation with low computational cost, though its reliance on frozen backbones may restrict fine-grained temporal reasoning compared to full or partial fine-tuning.
Discussion: The aforementioned image-to-video transfer learning methods primarily adapt existing image–text foundation models to address video–text understanding tasks through two main directions: enhancing the encoders of established video understanding frameworks or incorporating additional learnable temporal modules into image–language foundation models. Notably, these approaches are not mutually exclusive. Existing studies have shown that their integration can yield complementary benefits and further improve transfer performance. For example, fine-tuning with auxiliary models can be seamlessly combined with parameter-efficient strategies such as LoRA, facilitating the development of more advanced and robust video understanding models.
IV Fine-grained Video-Text Understanding
In the following, we review methods that adapt existing pre-trained image-text foundation models to address specific video-text understanding tasks. Based on the required granularity of understanding, we categorize the video-text understanding tasks into fine-grained and coarse-grained classes, providing a clearer structural understanding and enabling researchers to identify methods most relevant to their research domains. We first introduce fine-grained video-text understanding, which requires precise localization of spatial regions (e.g., bounding boxes or segmentations), temporal intervals, or targeted editing of relevant video segments based on textual descriptions.
IV-A Multi-Object Tracking
Open Vocabulary Multi-Object Tracking (OV-MOT) and Referring Multi-Object Tracking (RMOT) are two multi-object tracking tasks extended from traditional MOT, in which the objects of interest are explicitly referred with natural language descriptions. Specifically, OV-MOT aims to track all the objects  [^8] In this setting, the object number could be more than one and vary for different samples.described by input text in an open-vocabulary manner, rather than relying on a predefined, closed set of categories. The performance of OV-MOT systems is measured through metrics including localization accuracy (LocA), association accuracy (AssocA), and classification accuracy (ClsA). In contrast, Referring Multi-Object Tracking (RMOT) takes a natural language description specifying the objects of interest as input, and aims to detect and track these objects within a given video. This task is challenging as it requires the model to jointly capture fine-grained visual appearance, motion patterns, and semantic language cues for maintaining persistent trajectories of referred entities.
Most of existing OV-MOT methods [107, 41, 32, 29, 28, 40, 31, 30]mainly follow a frozen feature paradigm to transfer useful knowledge from pre-trained CLIP and Grounding-DINO models. For instance, OVTrack [28] and OVTrack+ [29] develop knowledge distillation blocks to enhance the RoI features extracted by ResNet50 and FPN with pre-trained CLIP, Based on OVTrack, VOVTrack [30] further introduce object appearance self-consistency and spatial-appearance mutual-consistency losses to capture spatial relationships across frames, where object appearance/attribute information are extracted with CLIP. SLAck [31] also distills textual knowledge from CLIP into its semantic head and introduces a spatial-temporal object attention mechanism to jointly model semantic, location, and appearance cues. GLATrack [32] extracts semantic information about the target objects from the reference frame using pre-trained CLIP model, which is then employed to reduce the influence of erroneous features in key frames. OVTR [41] employs a post-network fine-tuning strategy, wherein a Transformer-based dual-branch decoder is introduced to process frozen features extracted by CLIP. These methods rely on coarse-grained image-text correspondences captured by CLIP, which restricts the model’s ability to capture the open-world objects with high dynamicity. MASA [107] introduces a plug-and-play tracking adapter to enhance open-world foundation models (e.g., Grounding-DINO) for tracking arbitrary detected objects. NetTrack [40] introduced a plug-and-play framework that integrates a grounding model (e.g., Grounding-DINO[142]) to establish fine-grained object-text correspondences and a point tracker module (e.g., CoTracker[143]) to trace the trajectories of points of interest, thereby enhancing cross-frame object association performance.
Earlier RMOT approaches [144, 145, 146] mainly buit their systems based on a pretrained single-modal detection model (Deformable DETR), with cross-modal alignment and tracking modules integrated into this framework. Recently, iKUN [88] utilizes CLIP to encode the input query and each individual frame in video tubes, and subsequently employs a tracker to generate object trajectories. Furthermore, it introduces a knowledge unification module to guide the integration of visual features and textual information, producing text-specific visual representations and facilitating the selection of the most relevant tracking tubes. LaMOTer [42] conducts object tracking by detecting target objects in each frame with Grounding-DINO, predicting their locations in the next frame using an improved Kalman filter, and associating them to form the final tracks.
IV-B Video Segmentation
Open Vocabulary Video Instance Segmentation (OV-VIS) extends traditional VIS by enabling the segmentation and tracking of arbitrary object categories not seen during training, leveraging vision-language models to generalize beyond a fixed label set. In current research, the mainstream approach for this task leverages CLIP’s open-vocabulary capabilities by designing additional network modules appended after the CLIP encoder to achieve image-to-video transfer. OV2Seg [43] constitutes the inaugural work to incorporate CLIP into a post-network fine-tuning strategy for OVVIS. It employs a DETR-like architecture with object-centric queries to predict masks and perform cross-frame matching for tracking. The query embeddings, which fuse information from current and previous frames, are directly compared with CLIP-encoded text features to achieve open-vocabulary classification. InstFormer [97] introduces instance-guided attention into CLIP’s visual encoder, leveraging segmentation results as external cues to aggregate text-aligned instance features without compromising CLIP’s pretrained vision-language matching knowledge (using LoRA [141]). Furthermore, it feeds the aggregated instance tokens into the tracking module to match instance tokens between previous frames and the current frame. BriVIS [51] employs a frozen video instance segmentor to generate frame-level instance queries and integrates multi-scale CLIP visual features for alternating short-term and long-term post-network temporal modeling. To model instance dynamics, BriVIS introduces a Brownian bridge to constrain instance features and adopts a contrastive objective, ensuring that the bridge center remains close to its corresponding class text embedding while distancing itself from irrelevant class texts. OVFormer [52] employs video-level instance queries, as opposed to image-level queries, and aligns video-instance queries with frozen CLIP features via a cross-attention mechanism to achieve unified segmentation and tracking across frames. OV-DVIS++ [53] decomposes the task into segment-tracking-refinement stages, progressively modeling the problem from local to global contexts. CLIP-VIS [44] directly leverages CLIP’s image features to generate masks by combining predicted object scores and mask IoU. To alleviate the effect of historical noise, it first applies the Hungarian algorithm to identify the $K$ most relevant frames from memory, and then performs instance tracking.
Referring Video Object Segmentation (RVOS) aims to segment and track a target object instance across video frames based on given natural language description, which typically specifies the object’s appearance, motion, and relational cues. Grounded-SAM-2 [98] and AL-Ref-SAM-2 [99] follow the framework of finetuning with extra model, which attempt to employ GroundingDINO [142] to detect objects in individual frames, then use the grounding results as prompts for independent segmentation foundation models (e.g., SAM-2 [147]) to produce high-quality object masks. ReferDINO [72] proposes an end-to-end learning framework based on GroundingDINO, which first introduce a object-consistent temporal enhancer to capture the cross-frame object interaction. Then, it build a query pruning block on top of the pre-trained multimodality encoders in GroundingDINO, explicitly reducing the number of tokens for efficient inference. By combining the above modules and a mask decoder, ReferDINO succesfully empower the image-based foundation model with the ability of cross-modal temporal reasoning and segmentation.
IV-C Temporal Video Grounding
Temporal Video Grounding (TVG) is an important fine-grained multimodal video understanding task to localize the temporal boundaries of given language queries, which are typically in the form of single sentences [148, 149, 45], or even more complicated formats like paragraphs [150, 151, 152, 153] or articles [154, 155]. Early approaches in TVG predominantly relied on a two-stage and proposal-based pipeline, where a large number of predefined candidate temporal segments are generated and then ranked according to their relevance to the language query. Afterwards, the field has shifted towards methods which are in a one-stage pipeline and those that can leverage powerful pre-trained image foundation models like CLIP [1] via image-to-video transfer learning. A pivotal development was the introduction of DETR-style frameworks [156] that follow a Post-Network Tuning scheme. Moment-DETR [45] pioneered this by reformulating TVG as a direct set-prediction problem, utilizing a trainable Transformer-based head over frozen backbone encoders. Building on this, BAM-DETR[46] introduced a boundary-oriented formulation and a dual-pathway decoder to achieve more precise temporal alignment. UniVTG[33] proposed a unified multi-modal grounding head to handle multiple tasks within the same framework. UnLoc[47] employed a post-pretraining scheme to fuse pyramid features from a CLIP model for both frames and text.
Beyond the post-network paradigm, other transfer strategies have been explored. VDI [60] proposes a language-side adaptation paradigm for video-language models. While keeping the CLIP visual encoder frozen, it adapts the text encoder by infusing it with video-derived supervisory signals via a side-tuning approach. R2-Tuning [59] introduced a Side-Tuning method in the form of a reversed recurrent tuning block, which performs query-modulated spatial pooling and coarse-to-fine temporal refinement. To incorporate explicit relational reasoning, PaTF[61] encodes external cues by introducing a parallel transformer that processes structured scene graph representations.
More recently, the advent of Multimodal Large Language Models has introduced a new research direction focused on enhancing their temporal awareness. LLaViLo[108] was the first to utilize a video semantic modeling adapter to inject a joint feature representation into an LLM. To incorporate explicit time information as external cues, Chrono [96] introduced a sequence-to-sequence approach that integrates video frames with their corresponding timestamp cues. Similarly, VTG-LLM [100] integrated timestamp knowledge through sequence-time embeddings, absolute-time tokens, and slot-based token compression. ReVisionLLM [109] incorporated a recursive processing architecture with hierarchical adapters to progressively refine temporal segments in hour-long videos. In a training-free manner, TFVTG [48] employed LLMs to reason about sub-event relationships while using separate VLMs to localize each sub-event. NumPro [95] employs a fine-tuning strategy augmented with explicit visual cues. It overlays unique numerical identifiers onto each video frame, thereby translating the temporal grounding task into a visual OCR problem that can be directly resolved by a Video LLM.
IV-D Spatio-Temporal Video Grounding
Spatio-Temporal Video Grounding aims to predict a spatio-temporal tube for the event depicted in given language query from an untrimmed video. HCSTVG-v1, HCSTVG-v2, VidSTG-Declarative, and VidSTG-Interrogative are the most popular benchmarking datasets used in the community. The evaluation metrics for this task include: (1) m_vIoU, vIoU@0.3, and vIoU@0.5, which measure the spatiotemporal localization accuracy; and (2) m_tIoU, which solely evaluates the temporal localization precision.
Earlier methods [157, 158, 159] use unimodal detectors like Faster R-CNN [160] to detect objects in each video frame and then ground the temporal locations based on the detection features. STVGBert [68] is the first work to introduce and fully finetune a pre-trained ViLBERT [161] model and develop a novel spatio-temporal composition/decomposition attention module, thereby enabling joint spatio-temporal modeling. Recent studies [70, 71, 89, 162, 90, 92, 91] have developed frameworks based on the powerful pre-trained MDETR model [163], incorporating additional spatio-temporal interaction mechanisms. To achieve better performance, some approaches [70, 71] additionally introduce a temporal module to capture the cross-frame interaction and fine-tune all the parameters involved in the model. Specifically, TubeDETR [70] introduces a fast visual-only branch in the encoder and a time-aligned cross-attention mechanism in the decoder, enabling object queries to achieve joint spatio-temporal modeling. STCAT [71] designs a video-level multimodal template in the decoder to maintain temporal consistency across spatial predictions.
Beyond adapting image backbones for dynamic information modeling, recent SOTA methods further incorporate video-level features from additional video backbones to enhance the capture of temporal dynamics. CoSD [89] leverages SlowFast[137] features and proposes a two-stream framework with cross-stream collaboration to capture static appearance and dynamic motion cues collaboratively for target localization. CoSTA [90] employs a static-dynamic encoder-decoder architecture, while additionally utilizing video features extracted by 3D-CNN [164]. During the decoding process, it entangles temporal queries with spatial features and bridges temporal and spatial localizations. CGSTVG [91] and TA-STVG [92] integrate VidSwin [138] and MDETR to extract target-aware spatial and temporal cues under textual guidance.
In weakly-supervised STVG, the recent SOTA methods have increasingly adopted post-network tuning strategies. The statical vision-language alignment knowledge learned by pre-trained grounding models (like Grounding-DINO [142]) is explicitly exploited in the community, alleviating the requirement of manual spatiotemporal annotations in the training stage. For example, CoSPaL [49] and STPro [50] leverage pretrained Grounding-DINO to generate frame-wise spatial predictions, which are then concatenated to form a set of candidate tubelets with annotations. They then develop a curriculum learning strategy to capture spatio-temporal correspondences between sentence and tubelets, filtering out non-relevant proposals.
V Coarse-grained Video Understanding
Compared to fine-grained video-text understanding, coarse-grained tasks do not require precise temporal grounding results. They place greater emphasis on the holistic comprehension of events depicted in both text and video, such as determining whether a text-described event occurs in a video (Video Question Answering) or generating textual descriptions of video content (Video Captioning). In this section, we review methods for coarse-grained video-text understanding tasks from the perspective of transferring Image-Language Foundation Models to video domains.
V-A Video-Text Retrieval
Video–Text Retrieval (VTR) aims to retrieve the most relevant video from a large corpus given a textual query, or vice versa. It serves as a fundamental task in video understanding and often acts as a precursor to more fine-grained video–language tasks. Since the introduction of the large-scale image–text pretrained model CLIP [1], a series of subsequent studies have focused on transferring its knowledge to video–text retrieval, achieving substantial performance gains over earlier methods. Early approaches in this domain typically adopt a full fine-tuning paradigm, where model parameters are initialized from pretrained weights and subsequently optimized for downstream tasks. that explored transferring CLIP’s image representations to the video domain for video–text retrieval. It employs a 2D or 3D linear projection layer on flattened patches and introduces an additional learnable token to extract frame-level embeddings, following a pipeline similar to that of the vanilla ViT [165]. In addition, a preliminary video-text post-pretraining approach was explored in CLIP4Clip to further adapt CLIP representations to the video domain. CLIP2Video [63] extends image–text pretrained models to video–text retrieval by introducing a temporal difference block and a temporal alignment block. The temporal difference block enhances motion-related cues by modeling differences between adjacent frames, while the temporal alignment block aggregates text and frame embeddings through learnable center embeddings in the joint space. Another method [80] improves the full fine-tuning performance of CLIP for VTR by developing a weighted token-wise interaction mechanism and a channel de-correlation regularization loss. CLIP-Hitchhiker [78] observed that most single-modal temporal modeling and aggregation strategies under-perform a simple average temporal pooling operation. To address this, it introduced a query-guided temporal aggregation baseline, which achieved performances superior to state-of-the-art methods. CenterCLIP [81] proposes a more efficient full fine-tuning paradigm to transfer CLIP’s knowledge, where token clustering is conducted within each video segment and only center tokens are retained and further processed to compute the video-level embedding. TS2-Net [115] combines a temporal token shifting operation and a spatial token selection module to enhance the spatial-temporal video representation learning for better retrieval. CLIP2TV [35] constructs soft labels from the intra-modality similarity matrices to guide the learning of a fusion-based video-text matching module. X-CLIP [124] transfers the knowledge of pretrained CLIP models to end-to-end video–text retrieval through multi-grained contrastive learning across video–sentence and frame–word representations. In addition, an Attention over Similarity Matrix (AOSM) mechanism is employed to integrate similarity scores across multiple cross-modal levels. CLIP-ViP [126] introduces video proxy tokens as intermediaries for local–global information exchange, and further conducts video–text post-pretraining followed by downstream fine-tuning. To mitigate the injection of irrelevant visual information during video representation learning, X-pool [101] introduces a text-conditioned frame embedding pooling module, enabling more effective transfer of CLIP for video–text retrieval. VOP [125] introduces a cooperative text–video prompt-tuning strategy to mitigate catastrophic forgetting of pretrained vision–language knowledge while substantially reducing the number of trainable parameters. Considering the mutual dependency between contrastive dual encoders and the spatio-temporal characteristics of video data, VOP inserts learnable prompt tokens at the inputs of each text and vision encoder block, with positional, contextual, and functional aspects of video prompts explicitly modeled to enable effective transfer. CLIPPING [38] proposes an efficient knowledge distillation framework with a novel Student-As-Base (SAB) scheme, which regards the student’s intermediate features as the bases of the teacher’s feature space to absorb the pretrained knowledge. PromptSwitch [166] inserts a cube tensor as learnable prompts into the CLIP image encoder, where the prompt cube can be flexibly transposed to assist in the peer-to-peer communication between each frame pairs in the video. RAP [167] designs a low-rank modulation operation in the adapter module to inject the temporal sparsity and correlation characteristics for efficient image-to-video transfer. MV-Adapter [168] equips the adapter module with two novel components of temporal adaptation and cross modality typing that can effectively tackle the task of parameter-efficient video-text retrieval. TeachCLIP [39] introduces a multi-grained knowledge distillation framework that leverages a frozen CLIP model as the teacher. It jointly performs fine-grained frame-level distillation and coarse-grained video-level distillation, leading to strong performance in video–text retrieval. DiscoVLA [119] identifies three key image-to-video transfer gaps including vision, language and alignment and then introduces the IVFusion and PImgAlign to accordingly address the three transfer gaps for parameter-efficient video-text retrieval.
V-B Video Action Recognition
Video action recognition is a classical task in video understanding, which aims to recognize the category of action being performed in trimmed video. Over the years, there have been plentiful research works making efforts to transfer the powerful pretrained image-text foundation models for tackling video action recognition. ActionCLIP [121] is a pioneering work that adopts a “pretrain, prompt, and fine-tune” paradigm for transferring the knowledge of CLIP, which formulates action recognition into a multimodal learning problem by extracting text representations of class labels for video-text matching. X-CLIP [116] introduces a new framework to expand existing large-scale image-text pretrained models for general video recognition, where a cross-frame communication attention module and a video-specific textual prompting strategy are incorporated for better transfer performance. EVL [65] proposes to utilize a sideway transformer decoder to progressively aggregate multiple intermediate features from a frozen pretrained backbone to a learnable query token that is regarded as the global feature for predicting action classes. ST-Adapter [111] presents an efficient spatio-temporal adapter which is instantiated as a 3D convolution module inserted at the front end of each pretrained block in cascade. Adaptformer [112] designs another kind of adapter module which is inserted as a parallel learnable branch for merging into the original MLP’s output. AIM [113] jointly incorporates cascaded adapters into the spatial and temporal attention layers and adds an adapter in parallel to the MLP layer to achieve efficient transfer learning. BIKE [106] explores a bidirectional cross-modal knowledge transfer approach, where the attributes-category association and video concept spotting mechanisms are introduced to complement the unidirectional video-to-text matching and enhance the temporal awareness of video representations. ViFi-CLIP [79] shows that simple full fine-tuning of pretrained CLIP models with late temporal pooling generalizes well to supervised settings and a vision-language prompting scheme works well in low-data regimes. Text4Vis [36] demonstrates that a new paradigm with frozen classifiers initialized from pretrained textual embeddings can achieve more efficient transfer learning as compared to previous contrastive-based paradigms. DiST [66] disentangles spatial and temporal modeling to the pretrained backbone and a learnable network, and further utilizes a side branch to integrate multiple intermediate features from the spatial and temporal streams. UniFormerV2 [7] designs a sophisticated transfer method by additionally learning local temporal modeling and global spatio-temporal modeling on top of the pretrained global spatial modeling and a multi-stage fusion block is utilized to aggregate video tokens from different semantic levels. DUALPATH [114] explicitly separates the image-to-video adaptation into a spatial path with parallel adapters and a temporal path with serial adapters followed by a late path fusion for action classification. In addition, a new manner to imitate spatial modeling to achieve efficient temporal modeling via token interactions in a grid-like frameset is also proposed. MoMa [102] proposes a new temporal modeling method called sequence modulation operation, which is achieved by efficient trainable mamba layers to inject fine-grained spatio-temporal information into intermediate features. D2ST [117] develops a dual-pathway adapter that enables disentangled adaptation of spatial and temporal features, while maintaining full compatibility with pretrained single-stream image-text models.
V-C Video Captioning
Video captioning aims to generate natural language descriptions of video content. It has evolved from early template-based methods [169, 170, 171] to the adaptation of large pre-trained foundation models. Many works adopt a post-network tuning strategy to enhance features from frozen image encoders. Clip4Caption [56] introduces a post-network for temporal segment sampling and utilizes contrastive pretraining to learn text-aligned video embeddings from CLIP features. EvCap [105] also employs a post-network approach to fuse global visual contexts with specific action and object cues extracted by a pretrained CLIP model. mPLUG-2 [87] introduces a partial pretraining paradigm where a dual-vision encoder decouples spatial and temporal modeling by augmenting shared spatial self-attention layers with a novel local temporal modeling module for video inputs. MAMS [67] introduces a model-agnostic Side-Tuning framework that utilizes a module selector to adaptively choose between captioning modules of different sizes and incorporates an adaptive attention masking scheme to focus on important visual tokens. Similarly, CM2 [104] enhances CLIP features using multi-scale temporal convolutions and a cross-modal memory module that retrieves relevant textual semantics as external cues. In contrast, SWINBERT [69] proposes an end-to-end fine-tunable architecture, employing a Video Swin Transformer to extract dense spatio-temporal tokens which are then processed by a multimodal transformer with sparse attention for effective long-range modeling. More recent approaches focus on improving temporal precision. Vid2Seq [103] presents a unified framework for joint event captioning and localization by augmenting the language model’s vocabulary with special time tokens that serve as explicit temporal models. VTG-LLM [100] enhances pretrained encoders by injecting temporal cues like sequential and absolute time embeddings and employs a slot-based token compression mechanism for the efficient handling of long video sequences. HierarQ [58] leverages a trainable hierarchical Q-Former as post network, equipped with a two-stream feature modulator and dedicated memory banks, to compress frozen video features into a compact representation for a LoRA-adapted LLM.
V-D Video Question Answering
Pioneering methods [172, 173, 174, 175] either apply hierarchical spatio-temporal modeling strategies, employing two-stage architectures to first extract key frame clips followed by reasoning [176], or constructing spatio-temporal relational graphs via graph neural networks. Recently, more and more methods [120, 110] consider addressing the VideoQA problem utilizing the pre-trained foundation models.
Many studies have focused on exploring frozen-feature strategies, such as knowledge distillation and post-network tuning, for image-to-video transfer learning. VideoDistill [34] introduces a language-aware gating module that dynamically selects and refines frame-level visual features. This design facilitates goal-oriented selective knowledge distillation and effectively mitigates language bias that often arises when adapting image-based models to video data. BIMBA [54] adopts the post-network tuning strategy by introducing a Mamba-based token compression module on top of pre-trained foundation models. The module models long-range spatiotemporal dependencies and compresses redundant tokens, thereby improving efficiency. LinVT[55] implements post-network tuning through a plug-and-play Linear Video Tokenizer module. Placed after a frozen visual encoder, this post-network employs a Spatio-Temporal Visual Token Refiner (SVR) and a Text-Conditioned Token Aggregator (TTA) to produce a compact, question-relevant video representation for the subsequent large language model. HierarQ [58] develops a post-network tuning module in the form of a trainable hierarchical Q-Former equipped with task-aware modulators and memory mechanisms. It generates a compact yet informative spatiotemporal representation, which is subsequently fed into a frozen large language model that is further adapted through a lightweight adapter for textual answer generation. TEA [86] extends TextVQA [177] from images to videos by introducing spatio-temporal recovery modules into a pretrained T5 language model. These modules, consisting of a temporal convolution block and an OCR-enhanced relative spatial bias, can capture spatio-temporal clues to support video question answering.
Recent progress in Multimodal Large Language Models has established a new paradigm for Video QA, heavily reliant on the integration of modified visual and textual features. Atemporal Probe (ATP) [120] introduces a fine-tuning with prompt tuning approach that selects a single representative key frame to serve as a learnable visual prompt for a frozen vision encoder. Q-ViD [123] leverages a pre-trained InstructBLIP model [178] and incorporates tailored instructional prompts to generate descriptive video captions, using them as contextual cues for improved question answering. Similarly, ViTiS [122] employs multimodal prompt learning together with a visual mapping network to inject task-specific cues into image-language models, enabling strong performance with only few additional learnable parameters and limited training data. Other works explore adapter-based fine-tuning. Tem-Adapter [110] introduces a visual temporal aligner to predict future video states and a textual semantic aligner to refine embeddings for cross-modal alignment. TR-Adapter [118] proposes a temporal reasoning adapter designed to capture temporal relationships and facilitate knowledge transfer from language models. Unlike prior methods that maintain the original backbone architecture and perform adaptation through external lightweight modules, TC-LLaVA [84] modifies the attention computation in foundation models through partial tuning. Specifically, it introduces a Temporal-Aware Dual RoPE to encode temporal positional information and a Frame-wise Block Causal Attention Mask to enhance interactions among visual tokens.
[…]
V-E General Video Understanding
Unlike most existing image-to-video transfer learning methods which are tailored for one specific task, there have been some methods that aim to transfer the pretrained image-text knowledge across multiple tasks to achieve general video understanding. FitCLIP [37] proposes to train a student model by a small number of labeled video-text pairs and a large number of video-text pseudo labels, where the student model is then fused with the teacher model by weight ensembling for zero-shot video understanding. STAN [64] presents spatial-temporal auxiliary network as a side-tuning approach to perform temporal interaction and feature merging in the additional side branch, which achieves state-of-the-art performances across video-text retrieval and video action recognition.
TABLE III: Experimental results of representative methods with varied image-language foundation models and different image-to-video transfer learning strategies in coarse-grained video understanding tasks. Coarse-grained Video Understanding Task Category Method Venue ILFM R@1 $\uparrow$ R@5 $\uparrow$ R@10 $\uparrow$ R@sum $\uparrow$ VTR Full Fine-Tuning CLIP4Clip [77] Neurocomp. 2022 CLIP 44.5 71.4 81.6 - Finetune with Prompt Tuning VOP [125] CVPR 2023 CLIP 44.6 69.9 80.3 - Finetune with Adapter MV-Adapter [168] CVPR 2024 CLIP 46.2 73.2 82.7 - Knowledge Distillation TeachCLIP [39] CVPR 2024 CLIP 46.8 74.3 - 203.7 Finetune with LoRA DiscoVLA [119] CVPR 2025 CLIP 50.5 75.6 83.8 209.9 Task Category Method Publication ILFM K400 Top-1 $\uparrow$ K400 Top-5 $\uparrow$ SSv2 Top-1 $\uparrow$ SSv2 Top-5 $\uparrow$ VAR Side-Tuning EVL [65] ECCV 2022 CLIP 87.3 - 66.7 - Finetune with Adapter AIM [113] ICLR 2023 CLIP 87.5 97.7 70.6 92.7 Finetune with Adapter DUALPATH [114] CVPR 2023 CLIP 87.7 97.8 71.4 93.4 Side-Tuning DiST [66] ICCV 2023 CLIP 88.0 97.9 73.1 93.2 Finetune with Adapter ZeroI2V [179] ECCV 2024 CLIP 87.2 97.6 72.2 93.0 Finetune with Adapter MoMa [102] ICML 2025 CLIP 87.8 98.0 73.8 93.6 Task Category Method Publication ILFM MSRVTT-MC $\uparrow$ NextQA $\uparrow$ MSVD-QA $\uparrow$ How2QA $\uparrow$ VideoQA Finetune with prompt tuning ATP[120] CVPR 2022 CLIP 93.2 54.3 - 65.1 Finetune with Adapter Tem-Adapter[110] ICCV 2023 CLIP 94.3 - - - Finetune with prompt tuning ViTiS[122] ICCV 2023 CLIP - - 47.8 - Finetune with prompt tuning Q-Vid[123] ACL 2024 CLIP - 66.3 - 71.4 Knowledge Distillation VideoDistill[34] CVPR 2024 CLIP 97.8 - 49.2 - Post Network Tuning BIMBA[54] CVPR 2025 LLaVA - 83.7 - - Task Category Method Publication ILFM BLEU4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ Video Captioning Post Network Tuning CLIP4Caption[56] ACM MM 2021 CLIP 47.2 64.8 31.2 60.0 Full Fine-Tuning SWINBERT[69] CVPR 2022 CLIP 45.4 64.1 30.6 55.9 Finetune with partial tuning mPLUG-2[87] ICML 2023 CLIP 57.8 70.1 34.9 80.3 Fine-Tuning with Extra Models Vid2Seq[103] CVPR 2023 CLIP - - 30.8 64.6 Side-Tuning MAMS[67] AAAI 2025 - 60.0 71.2 34.7 82.9 Post Network Tuning HierarQ[58] CVPR 2025 LLaVA - - 35.1 80.5
VI Experimental Analysis
In the following, we present the experimental results of various transfer learning techniques applied to different video-text understanding tasks, followed by a detailed analysis.
VI-A Temporal Video Grounding
Datasets and Metrics. Temporal Video Grounding (TVG) aims to locate segments (or moment) in given untrimmed video based on a descriptive sentence. To address the TVG problem, numerous datasets have been constructed with video–text pairs and detailed annotations specifying the temporal spans of events described in the text. Specifically, DiDeMo[149], TACoS[180], Charades-STA[148] and ActivityNet Captions[181] are three most widely used benchmarks in the community. Following [148], most existing approaches are evaluated in these datasets and report their performance in the term of metrics ”Rn@m”and ”mIoU”. Metric ”Rn@m” is employed to compute the the percentage of samples where the top-$n$ predicted segments has an IoU greater than a predefined threshold $m$, where $m$ is typically defined as $\{0.1,0.3,0.5,0.7\}$. Metric ”mIoU” calculates the mean IoU between the predicted segments and ground truth on all testing data.
Results and Analysis. As shown in Table III, results for Temporal Video Grounding (TVG) reveal a notable performance gap between specialized architectures built on CLIP and generalist Multimodal Large Language Models (MLLMs) adapted from LLaVA. Among CLIP-based models, the Side-Tuning method R2-tuning achieves the highest precision with 59.8% on R@0.5 and 50.9% on m_I0U. This can be attributed to its ability to refine multi-level visual features within a dedicated temporal modeling module. In contrast, for MLLM-based methods, NumPro leads with 42.0% on R@0.5 and 41.4% on m_IoU by employing a Fine-Tuning with Extra Models strategy that reformulates temporal regression into a visual OCR task, better aligning with the MLLM’s native strengths. The observed results suggest that for TVG—a task requiring precise temporal regression—specialized, decoupled architectures currently hold an advantage over temporally-augmented general-purpose MLLMs.
VI-B Video-Text Retrieval
Datasets and Metrics. Video Text Retrieval (VTR) aims to identify relevant videos from a large collection based on a given descriptive sentence, or vice versa. To facilitate research in this area, several datasets have been constructed, pairing videos with textual descriptions and enabling the evaluation of cross-modal alignment performance. Among these, MSR-VTT [182] is one of the most widely used benchmarks in the field. Following common practice, most existing approaches are evaluated on this dataset and report performance using the metrics R@1, R@5, R@10, and R@sum. These metrics are defined as follows: R1, R5, and R10 measure the recall rate at the top 1, 5, and 10 retrieved results, respectively, indicating the model’s ability to rank relevant items at the top. Rsum is the sum of R1, R5, and R10, providing an aggregated measure of retrieval performance across different ranks.
Results and analysis. By examining the results presented in Table II for the task of Video-Text Retrieval, it can be observed that most recent methods are built upon the pretrained CLIP model due to its strong visual-textual correspondence capability. Among these, fine-tuning with LoRA achieves the best performance, significantly surpassing other feature-modification transfer learning strategies such as prompt tuning, knowledge distillation, and adapter-based tuning, in the metric of R@1. As expected, directly fusing CLIP-extracted features with a post-network tuning scheme performs worse than most of the approaches that updates CLIP parameters, highlighting the importance of adapting the original foundation model to the target task and dataset for optimal performance.
VI-C Video Action Recognition
Datasets and Metrics. Video Action Recognition (VAR) is a classical task in video understanding, which aims to recognize the category of human action being performed in a trimmed video. To support research in this area, several influential datasets have been developed. Foundational benchmarks include UCF101 [183] and HMDB51 [184], which established early standards for the task. More recently, the large-scale Kinetics series(e.g., Kinetics-400 [185], Kinetics-600 [186], and Kinetics-700 [187]) has become the most widely adopted benchmark for evaluating modern, high-capacity models. Following common practice, model performance on these datasets is primarily evaluated using Top-1 and Top-5 accuracy. Top-1 accuracy measures the percentage of videos for which the model’s single highest-probability prediction is correct. Top-5 accuracy measures the percentage of videos where the ground-truth label is among the model’s five most confident predictions, which is particularly informative for datasets with a large number of classes.
Results and analysis. Table II reports a detailed comparison of several representative image-to-video transfer methods for video action recognition based on the CLIP model. Overall, finetuning with Adapter strategy achieves better results than side-tuning strategy in this task. MoMa [102] obtains the best results on both K400 and SSv2 sets, which employs a mamba architecture to learn long-term fine-grained spatio-temporal information. The comparison between EVL and DiST demonstrates the effectiveness of disentangling the spatial and temporal modeling in image-to-video transfer learning.
VI-D Video Question Answering
Datasets and Metrics. Video Question Answering (VideoQA) task is typically evaluated using metrics such as Top-1 Accuracy, Top-K Accuracy, and Average Normalized Levenshtein Similarity (ANLS)[188], which measure the alignment between the predicted answers and ground truth answers. The VideoQA task is often subdivided into three distinct subcategories based on the nature of the questions and the expected answers: i):Multiple-Choice (MC): This subcategory involves answering questions that have a predefined set of possible answers. Common benchmarks for evaluating MC-based VideoQA tasks include TGIF-Action and TGIF-Transition[172], which focus on action-based queries within video clips, and MSRVTT-MC [182] and LSMDC-MC [189], which extend this challenge to larger, more diverse video datasets. ii):Open-Ended (OE): In this subcategory, the questions posed are more open-ended, allowing for subjective, creative, and sometimes logical responses. Popular benchmarks in the OE VideoQA subtask include TGIF-Frame [172], which deals with frame-level visual understanding; MSRVTT-QA [182], which presents questions based on the MSR-VTT video dataset; MSVD-QA [190], which uses the MSVD dataset for question answering; and LSMDC-FiB [189] and ActivityNet-QA [191], which extend the challenge to more diverse and dynamic video sequences. iii):Long-Form (LF): The Long-Form VideoQA[192] task takes VideoQA a step further by not only requiring a single answer but also necessitating the generation of comprehensive, detailed explanations. Key benchmarks in this subtask include ActivityNet-QA [191], How2QA [193], and VIOLIN [194], all of which emphasize the generation of detailed, explanatory answers that require a deep understanding of video context and reasoning across time.
Results and analysis. We present the comparison results on four datasets in Table II, where most competitors use CLIP as their backbone to ensure a fair comparison. For standard vision-language models such as CLIP, various transfer methods demonstrate strong and dataset-dependent performance. The distillation-based approach, VideoDistill [34], achieves state-of-the-art results on MSRVTT-MC and MSVD-QA with accuracies of 97.8% and 49.2%, respectively. In contrast, on the How2QA benchmark, the prompt-tuning method Q-Vid [123] attains the highest score at 71.4%. These results suggest that for a fixed backbone like CLIP, the most effective adaptation strategy varies across different benchmarks.
Conversely,the integration of Large Language Models as the foundation model substantially elevates performance. On the NextQA benchmark, the LLaVA-based BIMBA [54] model, which utilizes a post-network tuning strategy, achieves an accuracy of 83.73%. This represents a significant margin of improvement over the best CLIP-based method, Q-Vid with an accuracy of 66.3% on the same dataset. This finding underscores that post-network tuning is a potent transfer strategy for LLM-based models, as it efficiently adapts visual information to unlock the LLM’s powerful, built-in semantic reasoning for video understanding.
VI-E Video Captioning
Datasets and Metrics. A variety of standard benchmarks are collected to train and evaluate video captioning models. Prominent open-domain datasets include MSRVTT [195], comprising 10,000 video clips across 20 categories, and MSVD dataset [196]. Both datasets provide multiple human-annotated reference captions per video, making them ideal for evaluating models’ general-purpose descriptive capabilities and serving as foundational benchmarks for assessing performance across diverse scenarios. For caption tasks that require specialized procedural understanding, datasets such as YouCook2 [197] are commonly used. This dataset contains long instructional videos with temporally segmented annotations, challenging models to generate fine-grained, step-by-step descriptions of procedural content.
The performance of video captioning models is typically evaluated by comparing machine-generated candidate captions with a set of human-written reference captions using a suite of automated metrics. These metrics primarily measure both lexical similarity and semantic alignment between the generated and reference captions. For instance, BLEU@N [198] is a precision-oriented metric that measures n-gram overlap to evaluate fluency, while the recall-oriented ROUGE-L [199] uses the Longest Common Subsequence (LCS) to assess content fidelity. Building on these, METEOR [200] achieves better agreement with human judgment by considering synonyms and word stems, providing a more semantically informed evaluation of generated captions. CIDEr [201], developed specifically for captioning, measures consensus by representing sentences as TF-IDF–weighted n-gram vectors, rewarding captions that capture the salient information shared across the reference set. These metrics provide a holistic evaluation of caption quality, capturing both linguistic accuracy and semantic fidelity.
Results and analysis. As shown in Table II, recent advancements in video captioning are driven by modified transfer strategies that outperform traditional fine-tuning. The side-tuning method MAMS [67] achieves state-of-the-art performance across multiple key metrics (60.0 BLEU4, 71.2 ROUGE-L, 82.9 CIDEr). This success can be primarily attributed to its architectural design; by training a separate, lightweight module while keeping the foundation model frozen, it effectively learns task-specific temporal dynamics without the risk of catastrophic forgetting. The strong performance of the LLaVA-based HierarQ [58], which excels by aligning visual features with an LLM’s powerful reasoning, further validates this trend. These results indicate that the most effective transfer strategies are those that intelligently integrate new knowledge while preserving the core capabilities of the pre-trained backbone.
VI-F Spatio-Temporal Video Grounding
Datasets and Metrics. For the STVG task, VidSTG is currently the most widely-used benchmark dataset. It contains 99,943 video-text pairs, including 44,808 declarative sentences and 55,135 interrogative sentences from 10,303 videos spanning 80 object categories. For clearer comparison, we evaluate different methods under the declarative sentences setting. As for metrics, we follow previous works[91, 92] to use mean vIoU as the main metric. vIoU is computed as: $vIoU=\frac{1}{\left|T_{u}\right|}\sum_{t\in T_{i}}IoU(\hat{b}_{t},b_{t}),$ where $T_{i}$ and $T_{u}$ indicate the intersection and union between the time intervals obtained from annotation and prediction, respectively. $\hat{b}_{t},b_{t}$ are the predicted and ground truth bounding box for the $t$-th frame, respectively. We average the vIoU score over all the samples to obtain mean vIoU (termed m_vIoU). We also report vIoU@R which indicates the proportion of samples with vIoU higher than $R$. We also follow TubeDETR[70] and report the tIoU evaluating the temporal grounding performances. It is defined as: $tIoU=\frac{\left|T_{i}\right|}{\left|T_{u}\right|},$ where $T_{gt}$ indicates the set of frames inside the ground truth time interval.
Results and analysis. In fully-supervised STVG tasks, nearly all existing approaches adopt MDETR[163] as the foundation model, as it provides essential visual grounding capabilities and can be fully adapted to video-level understanding through fine-tuning. However, as shown in Table III, performing temporal modeling solely on the image-level features extracted by MDETR leads to limited temporal localization accuracy. Consequently, recent methods [89, 90, 91, 92] have incorporated auxiliary models to extract video-level features, such as SlowFast[137] and 3D-CNN[164], which substantially improve temporal localization performance, as evidenced by the increase in tIoU from 48.1 in TubeDETR [70] to 52.1 in CoSTA [90]. These results suggest that integrating both image-level and video-level features enables the model to achieve more precise fine-grained spatio-temporal localization. However, incorporating auxiliary models also increases computational cost; for example, TASTVG[92] was trained on 32 NVIDIA A100 GPUs for 15 hours.
VI-G Open Vocabulary Video Instance Segmentation
Datasets and Metrics. We compare the experimental results of all OVVIS methods on the LV-VIS[43] validation and test sets. LV-VIS is a large-vocabulary video instance segmentation dataset comprising 4,828 real-world videos spanning 1,196 categories. The videos are partitioned into 3,083 for training, 837 for validation, and 908 for testing. The categories are divided into 641 base categories—inherited from frequent and common classes in LVIS[202]—and 555 novel categories that are disjoint from the base categories. We adopt the Average Precision (AP) as the evaluation metric, computed as the mean over multiple IoU thresholds from 0.5 to 0.95. To better assess open-vocabulary performance, we also report APn, which denotes the average AP exclusively on novel categories.
Results and analysis. As shown in Table III, Recent OVVIS methods have adopted the paradigm of post-network tuning for image-to-video transfer learning, incorporating CLIP to introduce open-vocabulary capabilities. Through a comparative analysis of several CLIP-based methods, we observe that CLIP-VIS[44] achieves the best performances. This can be primarily attributed to its introduction of a temporal Top-K enhanced matching strategy. Instead of matching only adjacent frames or all historical frames, this approach selects the top K most similar frames from the past for query association. This mechanism reduces interference from noisy frames and mitigates error propagation, thereby enhancing tracking stability. Furthermore, the method integrates classification scores with an objectness score and a mask IoU score. This fusion strengthens the consistency between classification and segmentation quality, improving robustness against challenges such as occlusion and complex scenes.
VI-H Open Vocabulary Multi-Object Tracking
Datasets and Metrics. For the OVMOT task, we adopt the TAO[203] dataset as benchmark and compare the performance of existing image-to-video transfer methods. The TAO dataset is a comprehensive video collection designed to provide rich visual and semantic annotations for diverse video analysis tasks. A key characteristic of TAO is its open-world nature, enabling the tracking of arbitrary objects rather than restricting to predefined categories. The dataset is partitioned into three subsets: training, validation, and testing, comprising 500, 988, and 1,419 videos respectively.
The OVMOT method is primarily evaluated using the TETA metric. TETA[204], an extension of HOTA[205], enhances its functionality to address challenges in long-tail scenarios. The TETA score comprises three components: a localization score (LocA), a classification score (ClsA), and an association score (AssocA). LocA quantifies the precision of bounding box localization, measuring the spatial alignment between predicted and ground-truth boxes, irrespective of classification or association accuracy. ClsA evaluates the correctness of object classification into predefined categories (e.g., ’car’, ’pedestrian’, ’bicycle’), independent of localization or association performance. AssocA assesses the robustness of identity preservation over time, reflecting the consistency of object-ID assignments across frames and directly capturing tracking and data association efficacy. The overall TETA score is computed as the arithmetic mean of these three sub-scores.
Results and analysis. In the OVMOT task, researchers explored multiple image-to-video transfer paradigms, including distillation, adapter modules, and post-network tuning, while also experimenting with two distinct foundation models: CLIP[1] and GroundingDINO [4]. As shown in Table III, no single method clearly outperforms the others or achieves superior performance across all metrics. Specifically, the CLIP-based GLATrack model [32], which performs image-to-video transfer via knowledge distillation, achieves the highest scores on TETA and AssocA, surpassing other competitors by more than 1%. In contrast, the GroundingDINO-based approaches MASA [107] and NetTrack [40] perform better on LocA and ClsA, respectively, highlighting that different methods excel in different evaluation criteria.
VII FUTURE DIRECTIONS AND DISCUSSION
This section presents the challenges and potential research directions related to transfer useful knowledge from pre-trained vision-language foundation model for video-text multi-modality understanding.
VII-A Challenges
A key challenge in transferring vision-language foundation models to the video domain lies in the limited and specialized nature of their capabilities. For instance, CLIP is primarily designed for image-text alignment, BLIP focuses on vision-language pre-training for image captioning and question answering, MDETR emphasizes grounding textual descriptions in object detection tasks, while GroundingDINO targets phrase grounding with strong detection performance. Despite their effectiveness within individual domains, no unified model has yet demonstrated robust and comprehensive capabilities across diverse tasks. As a result, different application scenarios in video generation, understanding, or editing often require selecting distinct foundation models, leading to fragmented solutions. Moreover, a standardized and effective strategy for fine-tuning these models on video tasks has yet to be established. Current research predominantly investigates diverse transfer paradigms in an exploratory fashion, highlighting both the immaturity of existing approaches and the experimental stage of this research direction.
VII-B Future Directions
As the transfer of image-language foundation models to video-language tasks continues to evolve, several promising research directions emerge. Below, we outline three key trends that are likely to shape future efforts in this field.
Unified Transfer Learning Paradigm for Multiple Tasks. A significant and promising future direction is the development of a unified transfer learning framework that enables a single ILFM to be effectively adapted for multiple video-language tasks simultaneously. Current methods often require designing distinct fine-tuning strategies or adding task-specific modules for each target task (e.g., action recognition, temporal grounding, video QA), which is inefficient and lacks generalization. Future research should focus on creating a more universal adaptation mechanism. This could involve exploring prompt-based learning, where task-specific instructions guide the model’s behavior, or designing parameter-efficient tuning methods (e.g., unified adapters, hypernetworks) that share most parameters across tasks. The ultimate goal is a single, unified model that can perform a wide array of video understanding tasks through a simple and consistent interface, moving beyond specific solutions toward a more general-purpose video intelligence system.
Collaboration of Multiple Foundation Models. Another promising avenue involves effectively integrating multiple pre-trained foundation models—each specializing in a different modality or capability—to tackle a single video-language task. For example, combining large language models (LLMs) with vision foundation models could enhance contextual reasoning and temporal understanding in complex video tasks. Research efforts could focus on fusion mechanisms, knowledge distillation, and cross-model attention to harness complementary strengths while mitigating computational costs.
Advanced Fusion Methods for Dual-Modality Integration. Despite progress, effectively aligning and fusing visual and linguistic features across spatial and temporal dimensions remains challenging. Future studies may investigate more dynamic and efficient fusion techniques, such as cross-modal transformers, graph-based alignment, or lightweight feature interaction modules. Improving the synergy between video and text modalities will be critical for tasks requiring fine-grained understanding, such as spatio-temporal grounding or long-form video reasoning.
These directions highlight the need for more efficient, integrative, and scalable approaches to bridge image-language models and video-language applications, paving the way for richer video understanding systems.
VII-C More Discussion
This survey primarily reviews the literature on transferring knowledge from image–text foundation models to advance video–text understanding. Indeed, recent studies have increasingly explored extending these foundation models with image-to-video transfer learning techniques for video editing and generation. A brief overview of related works is provided here, and readers are referred to [206] for a more comprehensive and systematic survey.
Methods. Video generation aims to create visually coherent and semantically meaningful video sequences from various inputs, such as text or images. The advent of vision–language foundation models such as Stable Diffusion (SD) [75] has stimulated extensive research on video generation, a task that demands effective modeling of temporal coherence and dynamic motion. Early works, like ModelScope Text-to-Video [85], modified the latent diffusion U-Net by adding 3D convolutions and temporal attention to the frozen spatial attention block, enabling direct text-to-video synthesis in the latent space. At the same time, commercial systems such as Runway Gen-2 [207] began to demonstrate advanced text-to-video capabilities, likely based on optimized spatiotemporal latent diffusion frameworks. A different approach was proposed by AnimateDiff [208], which introduced a lightweight Motion Module trained independently and injected into frozen U-Nets of pre-trained SD checkpoints using Motion LoRA. This design decouples the learning of appearance from that of motion, thereby improving the efficiency of animation. Subsequent works like VideoCrafter1 [83] further refined spatiotemporal U-Net architectures by partial tuning, often using causal 3D convolutions. For image-to-video generation, I2VGen-XL [76] employed a hierarchical latent diffusion framework that upsampled videos progressively across multiple stages. More recently, Stability AI released Stable Video Diffusion (SVD) [74], which extends and fully fine-tunes image-based Stable Diffusion models by incorporating temporal layers. This design enables the generation of short, temporally consistent video clips and even multi-view outputs.
Text-driven video editing aims to generate high-quality videos aligned with a target prompt while preserving the source video’s spatio-temporal structure. The main challenge is balancing flexible semantic edits with temporal consistency. Significant advancements in video editing stem from leveraging large-scale, pretrained vision-language foundation models, particularly Stable Diffusion (SD)[75]. This marks a departure from traditional video processing pipelines that handled frames individually, often leading to flickering artifacts and quality degradation. By adapting models initially trained on static images, researchers now explicitly address the temporal dimension inherent in video data, aiming for coherent edits across sequences.
Pioneering work Tune-A-Video [62] transferred pretrained Stable Diffusion to video tasks through side-tuning spatio-temporal attention modules. It introduced sparse-causal and temporal attention layers within the U-Net to improve consistency, but required costly per-video fine-tuning. To overcome this limitation, Text2Video-Zero [209] proposed a training-free approach. It integrated motion dynamics using cross-frame attention and latent warping, though performance dropped on complex motions due to reliance on implicit cues. Subsequent zero-shot methods explored richer conditioning strategies. Runway Gen-1 [73] combined CLIP embeddings and depth maps with a finetuned convolution–attention U-Net. Pix2Video [134] applied DDIM inversion and depth-guided denoising, preserving structure by injecting latent features from earlier frames without training. ControlNet [136] was adapted in ControlVideo [210], which combined cross-frame attention and interpolation for multimodal structural guidance in a training-free manner. Rerender A Video [211] improved temporal smoothness with sparse causal attention, latent warping, and optical-flow–based fusion in both latent and pixel spaces. More recent works emphasize explicit motion and correspondence modeling. TokenFlow [135] introduced a key-frame–centric global attention strategy. FLATTEN [212] propagated pixel correspondences via optical flow within a training-free temporal-guided attention. Building on this, FRESCO [82] refined flow-based methods with occlusion masks and combined cross-frame flow attention with intra-frame similarity. Together, these developments highlight a shift from implicit motion cues toward explicit modeling, enabling more faithful and temporally coherent video editing.
Evaluation Datasets and Metrics. Video generation and editing rely heavily on large-scale datasets and comprehensive evaluation protocols. Commonly used datasets include WebVid [213] for text–video pairs, LAION [214] for large-scale image–text pairs, Kinetics [6] for human actions, HowTo100M [215] for instructional videos, and DAVIS for video editing benchmarks. These datasets provide rich motion, semantic, and structural priors essential for training and evaluation. System performance is typically assessed using objective metrics such as Fréchet Video Distance (FVD) [216] for temporal consistency and perceptual quality, Fréchet Inception Distance (FID) [217] and Inception Score (IS) [218] for fidelity, and CLIP-based measures [219] for text–video alignment. For editing tasks, frame-wise accuracy (Fram-Acc) and temporal coherence (Tmp-Con) are further employed to evaluate spatial preservation and temporal stability. Beyond these automated measures, human evaluation remains indispensable for judging higher-level aspects such as narrative coherence, creativity, and artifact detection.
Analysis. Recent studies have shown that temporal coherence can be enhanced by integrating temporal attention into diffusion backbones such as Stable Diffusion, either through partial fine-tuning or by incorporating information from preceding frames in a training-free manner. Furthermore, leveraging auxiliary modalities such as depth maps and optical flow strengthens temporal modeling and enables finer-grained spatial understanding, thereby improving the quality of both video generation and editing.
However, with the growing availability of open-source video foundation models such as SVD [74] and WAN [220], recent studies have shifted toward training and fine-tuning video-based foundation models, which have demonstrated superior performance as compared with image-to-video transfer methods built upon image–language foundation models.
VIII CONCLUSION
This survey presents a comprehensive overview of recent advancements in image-to-video transfer learning within multimodal learning, which, to the best of our knowledge, is the first of its kind. We begin by outlining the necessary background, including fundamental concepts and an introduction to common image-language foundation models (ILFMs). Subsequently, we summarize nine distinct transfer paradigms employed for image-to-video transfer learning. For each video task, we categorize existing ILFM-based methods according to their technical perspectives and the paradigms used for transferring knowledge from ILFMs. In the experimental section, we detail the evaluation settings and present a fair comparison of results to analyze the performance of different transfer paradigms across various tasks. Finally, we discuss several prevailing challenges and suggest promising future research directions for image-to-video transfer learning.